# Tianxiang Chen USTC, China Alibaba Group, China

### 04 Mar 2024 v1

@@ 5 Conclusion
In this paper, we propose Mamba-in-Mamba (MiM-ISTD) structure for efficient ISTD. We uniformly divide the image into patches (akin to visual sentences) and further split each patch to multiple smaller sub-patches (visual words). We devise a pure Mamba-based MiM pyramid encoder that encompasses stacked MiM blocks. Each MiM block contains an outer Mamba block to process the sentence embeddings and an inner Mamba block to model the relation among word embeddings. The visual word embeddings are added to the visual sentence embedding after a linear projection. Compared to present SOTAs, our method can achieve more efficient modelling of both the local and global information. Experiments on two datasets have demonstrate the effectiveness of our method.

### 08 Mar 2024 v2

@@ 5 Conclusion
In this paper, we propose a Mamba-in-Mamba (MiM-ISTD) structure for efficient ISTD. We uniformly divide the image into patches as visual sentences, and further split each patch to multiple smaller sub-patches as visual words. We devise a pure Mamba-based MiM hierarchical encoder that encompasses stacked MiM blocks. Each MiM block contains an outer Mamba block to process the sentence embeddings and an inner Mamba block to model the relation among word embeddings. The visual word embeddings are added to the visual sentence embedding after a linear projection. Experiments show that our method can achieve more efficient modelling of both local and global information.